The average weight of three men is 160 lbs. The middle person weighs twice as much as the lightest, but 30 lbs less than the heaviest. What is the weight of the middle person?
Answer by htmentor(1343)

Middle person weighs twice as much as the lighter person: m = 2l -> l = m/2
Middle person weighs 30 lbs less than the heaviest: m = h - 30 -> h = m+30
The average weight = 160 = (l+m+h)/3 -> l+m+h = 480
Substitute the values for l and h above and solve for m:
m/2 + m + (m+30) = 480
5m/2 = 450
m = 900/5 = 180
So the middle person weighs 180 lb.
Check:
90 + 180 + (180+30) = 480, and the average is 160